# Bootstrapping Estimators based on the Block Maxima Method

### v1

@@ 1 Introduction
Let $(\bm{X}_{t})_{t\in\mathbb{Z}}$ with $\bm{X}_{t}=(X_{t1},\dots,X_{td})^{\top}$ denote a stationary time series in $\mathbb{R}^{d}$. For block size $r\in\mathbb{N}$, let $\bm{M}_{r}=\max_{t=1}^{r}\bm{X}_{t}=(\max_{t=1}^{r}X_{t1},\dots,\max_{t=1}^{r}X_{t1})^{\top}$ denote the componentwise maximum over $r$ successive observations. In view of a version of the classical extremal types theorem for multivariate time series, the distribution of $\bm{M}_{r}$ may be approximated by a multivariate extreme value distribution, for sufficiently large block size $r$ (see Section 2.1 for details). Both the distribution of $\bm{M}_{r}$ and parameters related to its potential weak limit distribution are common target parameters in extreme value statistics, with applications in finance, insurance or environmental statistics among others (Beirlant et al., , 2004; Katz et al., , 2002; Philip et al., , 2020).
@@
Suppose $\bm{X}_{1},\dots,\bm{X}_{n}$ is an observed stretch from $(\bm{X}_{t})_{t\in\mathbb{Z}}$. The classical block maxima method for estimating parameters associated with the law of $\bm{M}_{r}$ (or its limiting law for $r\to\infty$) consists of dividing the sampling period $\{1,\dots,n\}$ into $m$ disjoint blocks of length $r$ (for simplicity, we assume that $n=mr$), and of using the sample of (disjoint) block maxima $(\bm{M}_{1,r}^{({\operatorname{db}})},\dots,\bm{M}_{m,r}^{({\operatorname{db}})})$ as a starting point for statistical methods; here $\bm{M}^{({\operatorname{db}})}_{i,r}=\max_{t=(i-1)r+1,\dots,ir}\bm{X}_{t}$ denotes the $i$th disjoint block maximum. This classical approach can often be improved by instead considering the sample of sliding block maxima $(\bm{M}_{1,r}^{({\operatorname{sb}})},\dots,\bm{M}_{n-r+1,r}^{({\operatorname{sb}})})$ as a starting point; here $\bm{M}^{({\operatorname{sb}})}_{i,r}=\max_{t=i,\dots,i+r-1}\bm{X}_{t}$ denotes the block maximum of size $r$ starting at “day” $i$. Note that this sample also provides a stationary, but strongly auto-correlated sample from the law of $\bm{M}_{r}$. It was shown in Bücher and Segers, 2018a ; Zou et al., (2021) among others that estimators that explicitly or implicitly involve empirical means of the sliding block maxima sample typically have the same expectation but a slightly smaller variance than the respective counterparts based on the disjoint block maxima sample.
@@
It seems natural that a lower estimation variance offers the possibility of constructing smaller confidence intervals for a given confidence level. The main goal of this work is to find universal practical solutions for this heuristic based on suitable bootstrap approximations. A challenge consists of the fact that the large sample asymptotics of the sliding block maxima method depend on a blocking of blocks approach involving some intermediate blocking parameter converging to infinity at a well-balanced rate, whence standard bootstrap approaches for time series like the block bootstrap (Lahiri, , 2003) would depend on such an intermediate blocking parameter as well (which must be avoided due to the typically small effective sample size in extreme value analysis). However, as we will show, a novel approach based on what we call the circular (sliding) block maxima sample allows for consistent distributional approximations without relying on such an intermediate parameter sequence. At the same time, the method is computationally attractive because it avoids recalculation of any block maxima for the bootstrap sample.
@@
In general, various variants of the bootstrap have been routinely applied in extreme value statistics (see, for instance, Eastoe and Tawn, , 2008; Huser and Davison, , 2014). However, to the best of our knowledge, respective statistical theory is only available for the peak-over-threshold method (Peng and Qi, , 2008; Drees, , 2015; Davis et al., , 2018; Kulik and Soulier, , 2020; Jentsch and Kulik, , 2021; de Haan and Zhou, , 2024). Concerning the block maxima method, within the regime where the block size is treated as a parameter sequence $r=r_{n}$ converging to infinity, the only reference we are aware of is de Haan and Zhou, (2024), who derive an asymptotic expansion for the tail quantile process of bootstrapped disjoint block maxima. However, their results require an underlying i.i.d. sequence, they only concern the disjoint block maxima method and they do not eventually imply desirable classical consistency statements. While our main focus is on the sliding block maxima method, we establish consistency results for the disjoint block maxima method as a by-product, for serially dependent data.
@@
Finally, we also derive limit results for empirical means and estimators based on the circular (sliding) block maxima sample, and prove that they show the same favorable asymptotic behavior as the sliding block maxima counterparts. As we will discuss in the conclusion, the result may be of independent interest when dealing with non-stationary situations involving, for example, trends; note that the latter is a typical use case for climate extremes.
@@
The remaining parts of this paper are organized as follows: basic model assumptions and known results on the disjoint and sliding block maxima method are collected in Section 2. The sample of circular block maxima is introduced in Section 3 and supplemented by a central limit theorem. Respective bootstrap schemes are proposed in Section 4, where we also provide respective consistency statements. The previous high-level results are applied to a specific estimation problem for univariate heavy tailed time series in Section 5. Finite-sample results on a large-scale simulation study and a small case study are presented in Sections 6 and 7, respectively. Section 8 concludes, and all proofs and technical conditions are collected in Sections 9-11. Some additional theoretical results and simulation results are deferred to a sequence of appendices.
Throughout, weak convergence of sequences of random vectors/distributions is denoted by $\rightsquigarrow$. The code used for the simulation study is available on Github, see Staud, (2024).
@@ 2.1 Basic model assumptions
An extension of the classical extremal types theorem to strictly stationary time series (Leadbetter, , 1983) implies that, under suitable broad conditions, affinely standardized maxima extracted from a stationary time series converge to the generalized extreme value distribution (GEV). This was generalized to the multivariate case in Hsing, (1989), where the marginals are necessarily GEV-distributed. We make this an assumption, and additionally require the scaling sequences to exhibit some common regularity inspired by the max-domain of attraction condition in the i.i.d. case (de Haan and Ferreira, , 2006).
@@
Let $(\bm{X}_{t})_{t\in\mathbb{Z}}$ denote a strictly stationary time series in $\mathbb{R}^{d}$ with continuous margins. There exist sequences $(\bm{a}_{r})_{r}=(a_{r}^{(1)},\dots,a_{r}^{(d)}))_{r}\subset(0,\infty)^{d},(\bm{b}_{r})_{r}=(b_{r}^{(1)},\dots,b_{r}^{(d)})_{r}\subset\mathbb{R}^{d}$ and $\bm{\gamma}=(\gamma^{(1)},\dots,\gamma^{(d)})\in\mathbb{R}^{d}$, such that, for any $s>0$ and $j\in\{1,\dots,d\}$,
@@
where $G$ denotes a $d$-variate extreme-value distribution with marginal c.d.f.s $G_{\gamma^{(1)}},\dots,G_{\gamma^{(d)}}$ (with $G_{\gamma}(x)=\exp(-(1+\gamma x)^{-1/\gamma})\bm{1}(1+\gamma x>0)$ the cdf of the GEV($\gamma$)-distribution) and where division by the vector $\bm{a}_{r}$ is understood componentwise, that is, $\bm{Z}_{r}=(Z_{r}^{(1)},\dots,Z_{r}^{(d)})$ with $Z_{r}^{(j)}=\{\max(X_{1}^{(j)},\ldots,X_{r}^{(j)})-b_{r}^{(j)}\}/{a_{r}^{(j)}}$ for $j\in\{1,\dots,d\}$.
Note that (2.1) and (2.2) may for instance be deduced from Leadbetter’s $D(u_{n})$ condition, a domain-of-attraction condition on the associated i.i.d. sequence with stationary distribution equal to that of $\bm{X}_{0}$ and a weak requirement on the convergence of the c.d.f. of $\bm{Z}_{r},$ see Theorem 10.22 in Beirlant et al., (2004).
Throughout this paper, we assume to observe $\bm{X}_{1},\ldots,\bm{X}_{n}$, an excerpt from a $d$-dimensional time series $(\bm{X}_{t})_{t\in\mathbb{Z}}$ satisfying Condition 2.1. Block maxima will be formed with respect to a block size parameter $r=r_{n}$ converging to infinity such that $r=o(n)$. For establishing the subsequent limit results, the serial dependence of $(\bm{X}_{t})_{t}$ will additionally be controlled via the Rosenblatt mixing coefficient (Bradley, , 2005). For two sigma-fields $\mathcal{F}_{1},\mathcal{F}_{2}$ on a probability space $(\Omega,\mathcal{A},\mathbb{P})$, let $\alpha(\mathcal{F}_{1},\mathcal{F}_{2})=\sup_{A\in\mathcal{F}_{1},B\in\mathcal{F}_{2}}|\mathbb{P}(A\cap B)-\mathbb{P}(A)\mathbb{P}(B)|.$ For positive integer $p$, let $\alpha(p):=\alpha\big{(}\sigma((\bm{X}_{t})_{t\leq 0}),\sigma((\bm{X}_{t})_{t\geq p})\big{)}$, with $\sigma(\cdot)$ denoting the sigma-field generated by its argument.
@@
For the block size sequence $(r_{n})_{n}$ it holds that, as $n\to\infty$:
@@ 2.2 Empirical means of rescaled disjoint or sliding block maxima
A central theoretical ingredient for establishing weak limit results on estimators based on the block maxima method is the weak convergence of either the tail quantile process or of empirical means of (unobservable) rescaled block maxima. Respective results on the former can be found in Theorem 2.1 in Ferreira and de Haan, (2015), while the latter approach was taken in Theorems 2.6/B.1 in Bücher and Zanger, (2023), in the proof of Theorem 2.6 in Bücher and Segers, 2018a or in Theorem 2.4 in Zou et al., (2021). Throughout this paper, we also follow the latter approach, and for completeness, we summarize the essence of the respective results in a theorem and briefly summarize and discuss potential statistical applications.
[…]
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{mb}})}-P_{r}\big{)},\quad\mathbb{G}_{n,r}^{({\operatorname{mb}})}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{({\operatorname{mb}})}-P\big{)},$
where $\mathbb{P}_{n,r}^{({\operatorname{mb}})}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{Z}_{r,i}^{({\operatorname{mb}})}}$, $P_{r}=\mathbb{P}(\bm{Z}_{r}\in\cdot)$ and $P=\mathbb{P}(\bm{Z}\in\cdot)$. Finally, let
@@
and note that $G_{\xi}$ is a $2d$-variate extreme-value distribution; see Formula (3.8) and its discussion in Bücher and Staud, (2024) for further details.
@@
Under Conditions 2.1 and 2.2, for any finite set of real valued functions $h_{1},\dots,h_{q}$ satisfying the integrability Condition 10.1(a) with $\nu>2/\omega$ where $\omega$ is from Condition 2.2, we have, writing $\bm{h}=(h_{1},\dots,h_{q})^{\top}$,
$\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}\bm{h}=\big{(}\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}h_{1},\dots,\bar{\mathbb{G}}_{n,r}^{({\operatorname{mb}})}h_{q}\big{)}^{\top}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{({\operatorname{mb}})}),\qquad{\operatorname{mb}}\in\{{\operatorname{db}},{\operatorname{sb}}\},$
@@
$\big{(}\Sigma_{\bm{h}}^{({\operatorname{db}})}\big{)}_{j,j^{\prime}=1}^{q}=\operatorname{Cov}\big{(}h_{j}(\bm{Z}),h_{j^{\prime}}(\bm{Z})\big{)},\qquad\big{(}\Sigma_{\bm{h}}^{({\operatorname{sb}})}\big{)}_{j,j^{\prime}=1}^{q}=2\int_{0}^{1}\operatorname{Cov}\big{(}h_{j}(\bm{Z}_{1,\xi}),h_{j^{\prime}}(\bm{Z}_{2,\xi})\big{)}\,\mathrm{d}\xi,$ (2.4)
with $\bm{Z}\sim G$ from Condition 2.1 and $(\bm{Z}_{1,\xi},\bm{Z}_{2,\xi})\sim G_{\xi}$ from (2.3). Moreover, $\Sigma_{\bm{h}}^{({\operatorname{sb}})}\leq_{L}\Sigma_{\bm{h}}^{({\operatorname{db}})}$, where $\leq_{L}$ denotes the Loewner ordering. Finally, if the bias Condition 10.2 holds, then
$\mathbb{G}_{n,r}^{({\operatorname{mb}})}\bm{h}=\big{(}\mathbb{G}_{n,r}^{({\operatorname{mb}})}h_{1},\dots,\mathbb{G}_{n,r}^{({\operatorname{mb}})}h_{q}\big{)}^{\top}\rightsquigarrow\mathcal{N}_{q}(\bm{B}_{\bm{h}},\Sigma_{\bm{h}}^{({\operatorname{mb}})}).$
Proof.
The proof is a simplified version of our proof of Theorem 3.2 below, whence we omit further details. Under slightly different conditions, a proof can be found in Theorem B.1 in Bücher and Zanger, (2023) or Theorem 8.7 in Bücher and Staud, (2024). ∎
In statistical applications, one is typically interested in distributional approximations for the estimation error of general statistics depending on the observable block maxima $\bm{M}_{r,i}$ (here and in the following, we omit the upper index ‘${\operatorname{mb}}$’). As discussed in the following remark, such approximations can be deduced from Theorem 2.3.
@@
Suppose that $\hat{\bm{\theta}}_{n}={\bm{\varphi}}_{n}(\bm{M}_{r,1},\dots,\bm{M}_{r,n})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$ (with $p\in\mathbb{N}$), where ${\bm{\varphi}}_{n}$ is a measurable function on $(\mathbb{R}^{d})^{n}$ with values in $\mathbb{R}^{p}$. As discussed in Example 2.5 below, for many functions ${\bm{\varphi}}_{n}$ of practical interest there exists a positive integer $q$ and functions ${\bm{\psi}}_{n}\colon(\mathbb{R}^{d})^{n}\to\mathbb{R}^{q}$, $A\colon(0,\infty)^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{p\times q}$ and $B\colon\mathbb{R}^{d}\times(0,\infty)^{d}\to\mathbb{R}^{p}$ such that

### v2

@@ 1 Introduction
A common approach to statistically assess extreme events is the block maxima method. For block size $r\in\mathbb{N}$, for instance $r=365$ corresponding to a year of daily observations, its starting point is the vector $\bm{M}_{r}=\max_{t=1}^{r}\bm{X}_{t}=(\max_{t=1}^{r}X_{t1},\dots,\max_{t=1}^{r}X_{td})^{\top}$ of componentwise maxima over $r$ successive observations from a time series $(\bm{X}_{t})_{t\in\mathbb{Z}}$ in $\mathbb{R}^{d}$; with $\bm{X}_{t}=(X_{t1},\dots,X_{td})^{\top}$. In view of a version of the classical extremal types theorem for multivariate time series, the distribution of $\bm{M}_{r}$ may be approximated by a multivariate extreme value distribution, for sufficiently large block size $r$. Both the distribution of $\bm{M}_{r}$ and parameters related to its potential weak limit distribution are common target parameters in extreme value statistics, with applications in finance, insurance or environmental statistics among others (BeiGoeSegTeu04; Kat02). For instance, in extreme event attribution studies (e.g., for cumulative daily precipitation amounts), one is typically interested in specific large quantiles or exeedance probabilities of $M_{365}$ that are defined by a recent observed extreme event. These parameters are directly linked to return levels and return periods of the underlying time series, and allow for assessing the effect of global warming on the observed event, for instance in terms of probability ratios (Van17; philip2020protocol; Tra23).
@@
Formally, the statistical analysis starts from $\bm{X}_{1},\dots,\bm{X}_{n}$, an observed stretch from an (approximately) stationary time series $(\bm{X}_{t})_{t\in\mathbb{Z}}$. The classical block maxima method for estimating quantities associated with the law of $\bm{M}_{r}$ (or its limiting law for $r\to\infty$) consists of dividing the sampling period $\{1,\dots,n\}$ into $m$ disjoint blocks of length $r$ (for simplicity, we assume that $n=mr$), and of using the (disjoint) block maxima sample $(\bm{M}_{1,r}^{[{\operatorname{d}}]},\dots,\bm{M}_{m,r}^{[{\operatorname{d}}]})$ as a starting point for statistical methods; here $\bm{M}^{[{\operatorname{d}}]}_{i,r}=\max_{t=(i-1)r+1,\dots,ir}\bm{X}_{t}$ denotes the $i$th disjoint block maximum. This classical approach can often be improved by instead considering the sliding block maxima sample $(\bm{M}_{1,r}^{[{\operatorname{s}}]},\dots,\bm{M}_{n-r+1,r}^{[{\operatorname{s}}]})$ as a starting point; here $\bm{M}^{[{\operatorname{s}}]}_{i,r}=\max_{t=i,\dots,i+r-1}\bm{X}_{t}$ denotes the block maximum of size $r$ starting at “day” $i$. Note that this sample also provides an (approximately) stationary, but strongly auto-correlated sample from the law of $\bm{M}_{r}$. It was shown in BucSeg18-sl; ZouVolBuc21 among others that estimators that explicitly or implicitly involve empirical means of the sliding block maxima sample typically have the same expectation but a slightly smaller variance than the respective counterparts based on the disjoint block maxima sample.
@@
It seems natural that a lower estimation variance offers the possibility of constructing smaller confidence intervals for a given confidence level, thereby for instance allowing for more precise statements regarding the aforementioned extreme event attribution. The main goal of this work is to find universal practical solutions for this heuristic based on suitable bootstrap approximations. A challenge consists of the fact that the large sample asymptotics of the sliding block maxima method depend on a blocking of blocks approach involving some intermediate blocking parameter converging to infinity at a well-balanced rate. As a consequence, standard bootstrap approaches for time series like the block bootstrap (Lah03) would depend on such an intermediate “blocking of blocks” parameter as well (which must be avoided due to the typically small effective sample size in extreme value analysis), while naive approaches based on (blockwise) resampling of sliding block maxima will be inconsistent (see Fig. 1, Remark 4.3 below, and Section D in the supplement). As illustrated on the right-hand side of Fig. 1, a novel approach based on what we call the circular (sliding) block maxima sample allows for consistent distributional approximations without relying on such an intermediate parameter sequence. At the same time, the method is computationally attractive because it avoids recalculation of any block maxima for the bootstrap sample.
Figure 1: Empirical distribution of the estimation error and the bootstrap estimation error for 100-year return level estimation in Model 6.1 with $\gamma=-0.2,\beta=0.5,m=80,r=365$. Left: naive sliding approach. Right: circular bootstrap approach. See Section 4 and Figure 4 for further details.
@@
In general, various variants of the bootstrap have been routinely applied in extreme value statistics (see, for instance, EasTaw08; HusDav14). However, to the best of our knowledge, respective statistical theory is only available for the peak-over-threshold method (Pen08; Dre15; Dav18; KulSou20; Jen21; DehZho24). Concerning the block maxima method, within the regime where the block size is treated as a parameter sequence $r=r_{n}$ converging to infinity, the only reference we are aware of is DehZho24, who derive an asymptotic expansion for the tail quantile process of bootstrapped disjoint block maxima. However, their results require an underlying IID sequence, they only concern the disjoint block maxima method and they do not eventually imply desirable classical consistency statements. While our main focus is on the sliding block maxima method, we establish consistency results for the disjoint block maxima method as a by-product, for serially dependent data.
@@
Finally, we also derive limit results for empirical means and estimators based on the circular (sliding) block maxima sample, and prove that they show the same favorable asymptotic behavior as the sliding block maxima counterparts. As we will discuss in the conclusion, the result may be of independent interest when dealing with non-stationary situations involving, for example, trends, which is typical for climate extremes.
All methods in this paper are intrinsically linked to the block maxima method. In typical use cases of that method, target parameters directly depend on the block maxima distribution (and hence on $r$); it can thus be considered a very natural approach in this situation. When target parameters depend on the marginal distribution of $\bm{X}_{t}$ (e.g., high quantiles), the peak-over-threshold method is typically considered the more natural and useful approach. As such, both methods have their raison d’être and are most suitable for different applications; see also the discussion in Section 3.3 in buecher2020horse.
@@
The remaining parts of this paper are organized as follows: basic model assumptions and known results on the disjoint and sliding block maxima method are collected in Section 2. The sample of circular block maxima is introduced in Section 3 and supplemented by a central limit theorem. Appropriate bootstrap schemes are proposed in Section 4, where we also provide respective consistency statements. The previous high-level results are applied to a specific estimation problem for univariate heavy tailed time series in Section 5. Finite-sample results on a large-scale simulation study and two case studies are presented in Sections 6 and 7, respectively. Section 8 concludes. Some technical integrability and bias conditions are given in Section 9. Finally, all proofs, additional theoretical results and additional simulation results are provided in a supplementary material. The code used for the simulation study is available on Github, see Sta24. Throughout, weak convergence of sequences of random vectors/distributions is denoted by $\rightsquigarrow$.
@@ 2 Mathematical preliminaries
Throughout this section, we present basic model assumptions rooted in extreme value theory for stationary time series. We then summarize the most important aspects about the (disjoint and sliding) block maxima method, including a central limit theorem (CLT) for sample averages based on affinely standardized block maxima. We also present high level conditions under which the CLTs can be used as a basis for statistical inference and illustrate them with some examples.
@@ 2.1 Basic model assumptions
An extension of the classical extremal types theorem to strictly stationary time series (Lea83) implies that, under suitable broad conditions, affinely standardized maxima extracted from a stationary time series converge to the generalized extreme value distribution (GEV). This was generalized to the multivariate case in Hsi89, where the marginals are necessarily GEV-distributed. We make this an assumption, and additionally require the scaling sequences to exhibit some common regularity inspired by the max-domain of attraction condition in the IID case (deextreme).
@@
Let $(\bm{X}_{t})_{t\in\mathbb{Z}}$ denote a strictly stationary time series in $\mathbb{R}^{d}$ with continuous margins. There exist sequences $(\bm{a}_{r})_{r}=(a_{r}^{(1)},\dots,a_{r}^{(d)})_{r}\subset(0,\infty)^{d},(\bm{b}_{r})_{r}=(b_{r}^{(1)},\dots,b_{r}^{(d)})_{r}\subset\mathbb{R}^{d}$ and $\bm{\gamma}=(\gamma^{(1)},\dots,\gamma^{(d)})\in\mathbb{R}^{d}$, such that, for any $s>0$ and $j\in\{1,\dots,d\}$,
@@
where $G$ denotes the cumulative distribution function (CDF) of a $d$-variate extreme-value distribution with marginal CDFs $G_{\gamma^{(1)}},\dots,G_{\gamma^{(d)}}$ (with $G_{\gamma}(x)=\exp(-(1+\gamma x)^{-1/\gamma})\bm{1}(1+\gamma x>0)$ the CDF of the GEV($\gamma$)-distribution) and where division by the vector $\bm{a}_{r}$ is understood componentwise, that is, $\bm{Z}_{r}=(Z_{r}^{(1)},\dots,Z_{r}^{(d)})$ with $Z_{r}^{(j)}=\{\max(X_{1}^{(j)},\ldots,X_{r}^{(j)})-b_{r}^{(j)}\}/{a_{r}^{(j)}}$ for $j\in\{1,\dots,d\}$.
Note that (2.1) and (2.2) may for instance be deduced from Leadbetter’s $D(u_{n})$ condition, a domain-of-attraction condition on the associated IID sequence with stationary distribution equal to that of $\bm{X}_{0}$ and a weak requirement on the convergence of the CDF of $\bm{Z}_{r},$ see Theorem 10.22 in BeiGoeSegTeu04.
Throughout this paper, we assume to observe $\bm{X}_{1},\ldots,\bm{X}_{n}$, an excerpt from a $d$-dimensional time series $(\bm{X}_{t})_{t\in\mathbb{Z}}$ satisfying Condition 2.1. Block maxima will be formed with respect to a block size parameter $r=r_{n}$ converging to infinity such that $r=o(n)$. For establishing the subsequent limit results, the serial dependence of $(\bm{X}_{t})_{t}$ will additionally be controlled via the Rosenblatt mixing coefficient (Bra05). For two sigma-fields $\mathcal{F}_{1},\mathcal{F}_{2}$ on a probability space $(\Omega,\mathcal{A},\mathbb{P})$, let $\alpha(\mathcal{F}_{1},\mathcal{F}_{2})=\sup_{A\in\mathcal{F}_{1},B\in\mathcal{F}_{2}}|\mathbb{P}(A\cap B)-\mathbb{P}(A)\mathbb{P}(B)|.$ For positive integer $p$, let $\alpha(p):=\alpha\big{(}\sigma((\bm{X}_{t})_{t\leq 0}),\sigma((\bm{X}_{t})_{t\geq p})\big{)}$, with $\sigma(\cdot)$ denoting the sigma-field generated by its argument.
@@
The block size sequence $(r_{n})_{n}$ satisfies, as $n\to\infty$:
@@ 2.2 Empirical means of rescaled disjoint or sliding block maxima
A central theoretical ingredient for establishing weak limit results on estimators based on the block maxima method is the weak convergence of either the tail quantile process or of empirical means of (unobservable) rescaled block maxima. Respective results on the former can be found in Theorem 2.1 in FerDeh15, while the latter approach was taken in Theorems 2.6/B.1 in BucZan23, in the proof of Theorem 2.6 in BucSeg18-sl or in Theorem 2.4 in ZouVolBuc21. Throughout this paper, we also follow the latter approach, and for completeness, we summarize the essence in a theorem and briefly summarize and discuss potential statistical applications.
For $i\in\{1,\dots,n-r+1\}$, write $\bm{M}_{r,i}=\max(\bm{X}_{i},\dots,\bm{X}_{i+r-1})$, and define ¯$\overline{\mathcal{M}}_{n,r}^{[{\operatorname{d}}]}=(\bm{M}_{r,i}:i\in I_{n}^{\operatorname{d}})$ and ¯$\overline{\mathcal{M}}_{n,r}^{[{\operatorname{s}}]}=(\bm{M}_{r,i}:i\in I_{n}^{\operatorname{s}})$ as the (vanilla) disjoint and sliding block maxima samples, respectively, where $I_{n}^{\operatorname{d}}=\{(i-1)r+1:1\leq i\leq n/r\}$ and $I_{n}^{\operatorname{s}}=\{1,\dots,n-r+1\}$. We are interested in the associated empirical measures $n_{\operatorname{m}}^{-1}\sum_{i\in I_{n}^{\operatorname{m}}}\delta_{\bm{M}_{r,i}}$, or their versions based on rescaled block maxima $n_{\operatorname{m}}^{-1}\sum_{i\in I_{n}^{\operatorname{m}}}\delta_{(\bm{M}_{r,i}-\bm{b}_{r})/\bm{a}_{r}}$, with $n_{\operatorname{m}}=|I_{n}^{\operatorname{m}}|$ and $\delta_{\bm{z}}$ the Dirac-measure at $\bm{z}$. The fact that the sample size $n_{\operatorname{m}}$ depends on ${\operatorname{m}}$ is a notational nuisance which we subsequently resolve by the following asymptotically negligible modification: first, the disjoint block maxima sample may be identified with the sample $\mathcal{M}_{n,r}^{[{\operatorname{d}}]}=(\bm{M}_{r,1},\dots,\bm{M}_{r,1},\dots,\bm{M}_{r,n/r},\dots,\bm{M}_{r,n/r})$ of size $n$ containing each disjoint block maximum exactly $r$ times; note that the respective empirical measures of ¯$\overline{\mathcal{M}}_{n,r}^{[{\operatorname{d}}]}$ and $\mathcal{M}_{n,r}^{[{\operatorname{d}}]}$ are the same. Next, we define $\mathcal{M}_{n,r}^{[{\operatorname{s}}]}$ as the sliding block maxima sample of size $n$ calculated from the extended sample $(\bm{X}_{1},\dots,\bm{X}_{n},\bm{X}_{n+1},\dots,\bm{X}_{n+r-1})$. In the subsequent asymptotic results, this modification is asymptotically negligible since the first $n-r+1$ maxima in $\mathcal{M}_{n,r}^{[{\operatorname{s}}]}$ are exactly the maxima in ¯$\overline{\mathcal{M}}_{n,r}^{[{\operatorname{s}}]}$. These modifications allow to define $(\bm{M}_{r,1}^{[{\operatorname{m}}]},\dots,\bm{M}_{r,n}^{[{\operatorname{m}}]}):=\mathcal{M}_{n,r}^{[{\operatorname{m}}]}$, and using the notation $\bm{Z}_{r,i}^{[{\operatorname{m}}]}:=(\bm{M}_{r,i}^{[{\operatorname{m}}]}-\bm{b}_{r})/\bm{a}_{r}$, we define
$\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{[{\operatorname{m}}]}-P_{r}\big{)},\quad\mathbb{G}_{n,r}^{[{\operatorname{m}}]}=\sqrt{\frac{n}{r}}\big{(}\mathbb{P}_{n,r}^{[{\operatorname{m}}]}-P\big{)},$
where $\mathbb{P}_{n,r}^{[{\operatorname{m}}]}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{Z}_{r,i}^{[{\operatorname{m}}]}}$, $P_{r}=\mathbb{P}(\bm{Z}_{r}\in\cdot)$ and $P=\mathbb{P}(\bm{Z}\in\cdot)$. Finally, let
@@
and note that $G_{\xi}$ is a $2d$-variate extreme-value distribution; see Formula (3.8) and its discussion in BucSta24 for further details. The following result relies on some suitable integrability and bias conditions (Conditions 9.1 and 9.2), which are postponed to Section 9 in the appendix for presentation reasons.
@@
Under Conditions 2.1 and 2.2, for any finite set of real valued functions $h_{1},\dots,h_{q}$ satisfying the integrability Condition 9.1(a) with $\nu>2/\omega$ where $\omega$ is from Condition 2.2, we have, writing $\bm{h}=(h_{1},\dots,h_{q})^{\top}$,
$\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}\bm{h}=\big{(}\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}h_{1},\dots,\bar{\mathbb{G}}_{n,r}^{[{\operatorname{m}}]}h_{q}\big{)}^{\top}\rightsquigarrow\mathcal{N}_{q}(\bm{0},\Sigma_{\bm{h}}^{[{\operatorname{m}}]}),\qquad{\operatorname{m}}\in\{{\operatorname{d}},{\operatorname{s}}\},$
@@
$\big{(}\Sigma_{\bm{h}}^{[{\operatorname{d}}]}\big{)}_{j,j^{\prime}=1}^{q}=\operatorname{Cov}\big{(}h_{j}(\bm{Z}),h_{j^{\prime}}(\bm{Z})\big{)},\quad\big{(}\Sigma_{\bm{h}}^{[{\operatorname{s}}]}\big{)}_{j,j^{\prime}=1}^{q}=2\int_{0}^{1}\operatorname{Cov}\big{(}h_{j}(\bm{Z}_{1,\xi}),h_{j^{\prime}}(\bm{Z}_{2,\xi})\big{)}\,\mathrm{d}\xi,$ (2.4)
with $\bm{Z}\sim G$ from Condition 2.1 and $(\bm{Z}_{1,\xi},\bm{Z}_{2,\xi})\sim G_{\xi}$ from (2.3). Moreover, $\Sigma_{\bm{h}}^{[{\operatorname{s}}]}\leq_{L}\Sigma_{\bm{h}}^{[{\operatorname{d}}]}$, where $\leq_{L}$ denotes the Loewner ordering. Finally, if the bias Condition 9.2 holds, then
$\mathbb{G}_{n,r}^{[{\operatorname{m}}]}\bm{h}=\big{(}\mathbb{G}_{n,r}^{[{\operatorname{m}}]}h_{1},\dots,\mathbb{G}_{n,r}^{[{\operatorname{m}}]}h_{q}\big{)}^{\top}\rightsquigarrow\mathcal{N}_{q}(\bm{B}_{\bm{h}},\Sigma_{\bm{h}}^{[{\operatorname{m}}]}).$
In statistical applications, one is typically interested in distributional approximations for the estimation error of general statistics depending on the observable block maxima $\bm{M}_{r,i}$ (here and in the following, we omit the superscript ‘${\operatorname{m}}$’). As discussed in the following remark, such approximations can be deduced from Theorem 2.3.
@@
Suppose that $\hat{\bm{\theta}}_{n}={\bm{\varphi}}_{n}(\bm{M}_{r,1},\dots,\bm{M}_{r,n})$ is some estimator of interest for a target parameter ${\bm{\theta}}_{r}\in\mathbb{R}^{p}$ (with $p\in\mathbb{N}$), where ${\bm{\varphi}}_{n}$ is a measurable function on $(\mathbb{R}^{d})^{n}$ with values in $\mathbb{R}^{p}$. Asymptotic normality of $\hat{\bm{\theta}}_{n}$ can often be deduced from asymptotic normality of some possibly different function evaluated in the (unobservable) rescaled block maxima $\bm{Z}_{r,1},\dots,\bm{Z}_{r,n}$; see, for instance, Seg01; BucZan23; Oor23; BucSta24. In this remark, we provide some high-level conditions under which such a conclusion is possible.
As discussed in Example 2.5 below, for many functions ${\bm{\varphi}}_{n}$ of practical interest there exists a positive integer $q$ and functions ${\bm{\psi}}_{n}\colon(\mathbb{R}^{d})^{n}\to\mathbb{R}^{q}$, $A\colon(0,\infty)^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{p\times q}$ and $B\colon\mathbb{R}^{d}\times(0,\infty)^{d}\to\mathbb{R}^{p}$ such that